OpenStudy (anonymous):

A leaf is torn from a paper back novel.The sum of the numbers on the remaining pages is 15000.What are the page numbers on the torn leaf?

OpenStudy (shubhamsrg):

let page nos be x and x+1, sum of pages torn off = 2x+1 let there be n pages initially, sum of all numbers = (n)(n+1)/2 we are given (n)(n+1)/2 - 2x - 1 = 15000 -4x + n^2 + n =30002 n^2 + n -4x -30002=0 we only need integral solutions of that, hmm am getting n=173 and x=25 ..please confirm..